2x+y+12=10
3 ordered pairs possible answers for x and y! Thanks a lot!:-)

Answer by Alan3354(69443)

2x+y+12=10
3 ordered pairs possible answers for x and y!
---------
Solve for y
2x+y+12=10
2x+y = -2
y = -2x -2
========
Pick any values for x and find y.